The compound interest on a certain sum for two years is Rs. 618 whereas the simple interest on the same sum at the same rate for two years is Rs. 600. The ratio of interest per annum is 1. 3%2. 18%3. 6%4. 9%

Grade:10

1 Answers

Vikas TU
7 years ago
compound interest is A = P (1 + r/n) nt
Hence given,
618 =  P (1 + r) 2
and
600 = P*r*2/100
dividing both we get,
618/600 = (10(1+r)/r)^2
we get r= 3% approx.